OpenStudy (anonymous):

Find the product (2x^2+6)(3-x^2)

OpenStudy (anonymous):

use FOIL method: http://calculator.tutorvista.com/foil-calculator.html

OpenStudy (anonymous):

(2x^2+6)(3-x^2) 18 - 2x^4 x = -√3 & √3

OpenStudy (shamim):

(2x^2+6)(3-x^2) =6x^2-2x^4+18-6x^2 =18-2x^4